# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v23

@@ 3.12 Hate Speech Detection
A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [266]. A feature model and a data set105105105https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [124]Sandaruwan et al. [267] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques. A cyberbullying comment classification study for Sinhala was conducted by [268] where they used classical machine learning algorithms. The study by Hettiarachchi et al. [269] used classical machine learning methods to detect hate speech in Romanized Sinhala social media posts. While the basic idea is the same, they have avoided mentioning transliteration in their paper. The study by Samarasinghe et al. [270] proposed using CNNs for detecting hate speech in Sinhala text. Kariyawasam [271] proposed a machine learning approach for identifying toxic Sinhala language on social media. Guruge et al. [272] used an ensemble of Naïve Bayes, Support Vector Machine, XGBoost, MLP, and AdaBoost to detect hate speech in 49019 Tweets they collected from February of 2021 to April of 2021. Sandaruwan et al. [273] collected 3000 Sinhala comments corpus and conducted Multinomial Naïve Bayes to detect hate speech. Sheran [274] claims to have used machine learning to detect hate speech written in Sinhala or Singlish on social media. A deep learning approach for detecting hate speech in Sinhala tweets was explored by Munasinghe and Thayasivam [275]. The study by Shalinda and Munasinghe [276] utilized classical machine learning techniques such as Linear Support vector classifier, Random Forest Classification, SGD classifier, Logistic Regression, XGBoost classifier and multinomial Naive Bayes classifier on both Sinhala and Singlish (Romanized Sinhala) to identify hate words. They report that the SGD classifier over TF-IDF with uni-grams and bi-grams gives them the highest accuracy. Gamage et al. [277] conducted a comparative analysis on a number of embedding systems as well as classical frequency-based methods for Sinhala hate speech detection. The study by Fernando et al. [278] also claims to use machine learning and deep learning to detect hate speech in Sinhala. The study by Perera et al. [279] predicted Sinhala hate speech using user behaviour on twitter by applying ensemble Machine Learning to classify them by the five big personality traits. The follow-up study by Perera et al. [125] analyses Sinhala hate speech propagation on Twitter. They have also published a dataset106106106https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. Rajapaksha et al. [280] used deep learning to identify trending periods of hate topics on Twitter in Sinhala. Arachchi et al. [281] used a web-based tool to translate between Sinhala and English to detect inappropriate word usage in Sinhala. The study by Ehelepola [282] analyzed Sinhala text from social media and e-commerce sites for hate speech. The work by Dikwatta et al. [283] used supervised algorithms to detect Sinhala hate speech in text from image posts on the Internet (memes). Wickramaarachchi et al. [284] used LSTM on BART to compare the title and description of Sinhala youtube videos to their audio in order to determine if they contain hate speech. The study by Muthuthanthri and Smith [285] used an annotated dataset from Facebook with TF-IDF and fastText embedding features with SVM, CNN, RNN, LSTM as well as BERT and GPT2. They concluded that BERT obtains the best result for detecting hate speech in English-Sinhala code-mixed data.
@@
After the release of the SOLD dataset by Ranasinghe et al. [135] for the hate speech detection task in Sinhala, a number of subsequent works have been conducted using it. Fernando and Deng [286] used traditional machine learning algorithms to predict Sinhala hate speech. Ranasinghe and Zampieri [287] compared the results of pre-trained mT5 on Sinhala tweets from the SOLD dataset against results for German, Spanish, Hindi, Korean, and Marathi. The work by Bestgen [288] compared the result of classical machine learning algorithms on simple n-gram features for the hate speech detection task on the SOLD dataset for Sinhala against results for Assamese and Bengali. Ranasinghe et al. [289] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task107107107https://hasocfire.github.io/hasoc/2023/task1.html they introduced. In completing the task, the study by Narayan et al. [290] compared the effectiveness of a number of pre-trained deep learning models for Sinhala hate speech detection. They showed that XLM Roberta Base achieves a macro-f1 of 83.49% over the LSTM baseline of 75.30%. For the same task, Ojo et al. [291] used mBERT to identify hate and offensive content in Sinhala.
@@ 3.14 Word Sense Disambiguation
There have been multiple attempts to do word sense disambiguation (WSD) [296, 297, 298, 299, 300] for Sinhala. For this, Arukgoda et al. [172] have proposed a system named Aruth based on the Lesk Algorithm[301]. An online tool108108108http://aruth.herokuapp.com/, an API109109109https://bit.ly/3sJEYbS of the algorithm, and code along with data on github110110110https://github.com/jseanm1/aruthSWSD are available. For the same task, Marasinghe et al. [302] have proposed a system based on probabilistic modeling. A dialogue act recognition system which utilizes simple classification algorithms has been proposed by Palihakkara et al. [303]. A word sense disambiguation tool named Sinsense was introduced by Subasingha [304]. They used cross-lingual sense disambiguation where English sense disambiguation was used to obtain Sinhala sense disambiguation. However, neither their tool nor their full research is publicly available.
@@ 3.15 Text Summarizing
A deterministic process flow for automatic Sinhala text summarizing was proposed by Welgama [305]. The study by Wimalasuriya [306], which has the same name as the above work by Welgama [305], uses the graph-based TextRank algorithm for automatic Sinhala text summarizing. The use case of Sinhala Text summarization for government gazettes was explored by Jayawardane [307]. The study by Rathnayake et al. [308] compared the results of extractive and abstractive summarization on Sinhala textbooks. The study by Jahan and Wijesekara [309] compared the abstractive summarizing methods of TF-IDF and Text-Rank for Sinhala using ROUGE as the evaluation score. They concluded TF-IDF to be the superior choice. Hasan et al. [144] introduced the XL-sum111111111https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website112112112https://www.bbc.com/sinhala. The M3LS113113113https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al. [145] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website114114114https://www.bbc.com/sinhala. Further, they claim that MT5 [310] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set. The first Sinhala multi-document summarizing dataset, M2DS115115115https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al. [146]. The M2DS dataset consists of 23.5k Sinhala documents in 5.5k clusters with each cluster having a golden summary. Thier code and pre-trained models are also available on github116116116https://github.com/KushanMH/m2ds.
@@ 3.17 Phonological Tools
On the case of phonological layer, a report on Sinhala phonetics and phonology was published by Wasala and Gamage [312]Wickramasinghe et al. [313] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems. The Massively Multilingual Speech (MMS)117117117https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al. [138] has Sinhala data for the Spoken Language Identification (LID) task. The meta-study by Al-Fraihat et al. [314] compared the status of Sinhala speech recognition research against 17 other languages.
@@ 3.17.1 Text-to-Speech
Based on the earlier work by Weerasinghe et al. [315]Wasala et al. [316] have developed methods for Sinhala grapheme-to-phoneme conversion along with a set of rules for schwa epenthesis. This work was then extended by Nadungodage et al. [317]Weerasinghe et al. [318] developed a Sinhala text-to-speech system. However, it is not publicly accessible. They internally extended it to create a system capable of helping a mute person achieve synthesized real-time interactive voice communication in Sinhala [319]. A rule based approach for automatic segmentation of a small set of Sinhala text into syllables was proposed by Kumara et al. [320]. An ew prosodic phrasing method to help with Sinhala Text-to-Speech process was proposed by Bandara et al. [321, 322, 323]Sodimana et al. [324] proposed a text normalization methodology for Sinhala text-to-speech systems. Further, Sodimana et al. [325] formalized a step-by-step process for building text-to-speech voices for Sinhala. Both Jayamanna [326] and Mishangi [327] have created Sinhala document readers for visually impaired persons to be used on Android devices. An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al. [328]. The works by Lakmal et al. [329] and Senarathna et al. [330] adapted MaryTTS [331] to synthesize Sinhala speech. The study by Jayawardhana et al. [332] used Deep Voice [333] for Sinhala and English TTS. Gamage et al. [334] included a Sinhala text-to-speech module as one of the three modules present in their currency recognition system. The study by Madhusha et al. [335] claims to have created a mobile app with Sinhala Text-to-Speech and OCR to read books for visually impaired students.  Anuradha and Thelijjagoda [336] proposed a machine translation system to convert Sinhala and English Braille documents into voice. A separate group has done work on Sinhala text-to-speech systems independent to above [337].
@@ 3.17.2 Speech-to-Text
On the converse, Nadungodage et al. [338] have done a series of work on Sinhala speech recognition with special notice given to Sinhala being a resource poor language. This project divides its focus on: continuity [339], active learning [340], and speaker adaptation [341]. A Sinhala speech recognition for voice dialing which is speaker independent was proposed by Amarasingha and Gamini [342] and on the other end, a Sinhala speech recognition methodology for interactive voice response systems, which are accessed through mobile phones was proposed by Manamperi et al. [343]. A Sinhala speech to Unicode text converter for the disaster relief domain was proposed by Prasangini and Nagahamulla [344]Priyadarshani [345] proposes a method for speaker dependant speech recognition based on their previous work on: dynamic time warping for recognizing isolated Sinhala words [346], genetic algorithms [347], and syllable segmentation method utilizing acoustic envelopes [348]. The method proposed by Gunasekara and Meegama [349] utilizes an HMM model for Sinhala speech-to-text. A Sinhala speech recognizer supporting bi-directional conversion between Unicode Sinhala and phonetic English was proposed by Punchimudiyanse and Meegama [350]. The work by Karunanayake et al. [351] transfer learns CNNs for transcribing free-form Sinhala and Tamil speech data sets for the purpose of classification. Dilshan [352] conducted a study for the specific use case of transcribing number sequences in continuous Sinhala speech.  Gamage et al. [353] explored the use of combinational acoustic models such as Deep Neural Network - Hidden Markov Model (DNN-HMM) [354] and Subspace Gaussian Mixture Model (SGMM) [354] in Sinhala speech recognition. In the follow-up work, Gamage et al. [355] extended that work with end-to-end Lattice-Free Maximum Mutual Information (e2e LF-MMI) model [356] which is claimed to be a viable solution for low resource language speech recognition by Carmantini et al. [357]. However, it was shown that the new model slightly underperforms compared to the state-of-the-art result. Later, they conducted further development on their model in a follow-up work [358].  Karunathilaka et al. [359] explore Sinhala speech recognition using deep learning models such as: pre-trained DNN, DNN, TDNN, TDNN+LSTM. The first half of the study by Arafath [360] dealt with recognizing Sinhala speech using LSTMs. Gamage et al. [334] included a Sinhala speech recognition module as one of the three modules present in their currency recognition system. Time-delay neural network architectures (including multistream CNN architecture) were used for acoustic modeling of Sinhala Automatic Speech Recognition (ASR) by Warusawithana et al. [361]. They have used the Kaldi speech recognition toolkit [362] for training the ASR models. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [363], proposed a Sinhala speech recognition system. The study by Azir et al. [364] attempts to identify number sequences spoken in Sinhala. TacoSi introduced by Arachchige and Weerasinghe [365] is based on Tacotron [366] and has been evaluated with 10 human evaluators to determine its text-to-speech quality. Nanayakkara [367], Nanayakkara and Weerasinghe [368] used DeepSpeech118118118https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition. Gunarathne et al. [369] used an earlier version of the CMUSphinx toolkit119119119https://cmusphinx.github.io/ to transcribe Sinhala speech to text. The later work by Akesh and Meegama [370] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [371, 372] to automatically generate Sinhala subtitles from Speech. The work by [373] also used MFCC features but focused on Sinhala speech intonation recognition for the purpose of detecting speech impediments in young children.

### v24

@@ 3.12 Hate Speech Detection
A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [290]. A feature model and a data set119119119https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [128]Sandaruwan et al. [291] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques. A cyberbullying comment classification study for Sinhala was conducted by [292] where they used classical machine learning algorithms. The study by Hettiarachchi et al. [293] used classical machine learning methods to detect hate speech in Romanized Sinhala social media posts. While the basic idea is the same, they have avoided mentioning transliteration in their paper. The study by Samarasinghe et al. [294] proposed using CNNs for detecting hate speech in Sinhala text. Kariyawasam [295] proposed a machine learning approach for identifying toxic Sinhala language on social media. Guruge et al. [296] used an ensemble of Naïve Bayes, Support Vector Machine, XGBoost, MLP, and AdaBoost to detect hate speech in 49019 Tweets they collected from February of 2021 to April of 2021. Sandaruwan et al. [297] collected 3000 Sinhala comments corpus and conducted Multinomial Naïve Bayes to detect hate speech. Sheran [298] claims to have used machine learning to detect hate speech written in Sinhala or Singlish on social media. A deep learning approach for detecting hate speech in Sinhala tweets was explored by Munasinghe and Thayasivam [299]. The study by Shalinda and Munasinghe [300] utilized classical machine learning techniques such as Linear Support vector classifier, Random Forest Classification, SGD classifier, Logistic Regression, XGBoost classifier and multinomial Naive Bayes classifier on both Sinhala and Singlish (Romanized Sinhala) to identify hate words. They report that the SGD classifier over TF-IDF with uni-grams and bi-grams gives them the highest accuracy. Gamage et al. [301] conducted a comparative analysis on a number of embedding systems as well as classical frequency-based methods for Sinhala hate speech detection. The study by Fernando et al. [302] also claims to use machine learning and deep learning to detect hate speech in Sinhala. The study by Perera et al. [303] predicted Sinhala hate speech using user behaviour on twitter by applying ensemble Machine Learning to classify them by the five big personality traits. The follow-up study by Perera et al. [129] analyses Sinhala hate speech propagation on Twitter. They have also published a dataset120120120https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. Rajapaksha et al. [304] used deep learning to identify trending periods of hate topics on Twitter in Sinhala. Arachchi et al. [305] used a web-based tool to translate between Sinhala and English to detect inappropriate word usage in Sinhala. The study by Ehelepola [306] analyzed Sinhala text from social media and e-commerce sites for hate speech. The work by Dikwatta et al. [307] used supervised algorithms to detect Sinhala hate speech in text from image posts on the Internet (memes). Wickramaarachchi et al. [308] used LSTM on BART to compare the title and description of Sinhala youtube videos to their audio in order to determine if they contain hate speech. The study by Muthuthanthri and Smith [309] used an annotated dataset from Facebook with TF-IDF and fastText embedding features with SVM, CNN, RNN, LSTM as well as BERT and GPT2. They concluded that BERT obtains the best result for detecting hate speech in English-Sinhala code-mixed data. Chavinda and Thayasivam [172] have used Multilingual Large Language Models (MLLMs) with Dual Contrastive Learning (DCL) to detect Sinhala hate speech on Facebook121121121https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and Twitter122122122https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech data sets. They report TwHIN-BERT-base [310] model to have obtained the best results.
@@
After the release of the SOLD dataset by Ranasinghe et al. [140] for the hate speech detection task in Sinhala, a number of subsequent works have been conducted using it. Fernando and Deng [311] used traditional machine learning algorithms to predict Sinhala hate speech. Ranasinghe and Zampieri [312] compared the results of pre-trained mT5 on Sinhala tweets from the SOLD dataset against results for German, Spanish, Hindi, Korean, and Marathi. The work by Bestgen [313] compared the result of classical machine learning algorithms on simple n-gram features for the hate speech detection task on the SOLD dataset for Sinhala against results for Assamese and Bengali. Ranasinghe et al. [314] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task123123123https://hasocfire.github.io/hasoc/2023/task1.html they introduced. In completing the task, the study by Narayan et al. [315] compared the effectiveness of a number of pre-trained deep learning models for Sinhala hate speech detection. They showed that XLM Roberta Base achieves a macro-f1 of 83.49% over the LSTM baseline of 75.30%. For the same task, Ojo et al. [316] used mBERT to identify hate and offensive content in Sinhala. The work by Rostamkhani and Eetemadi [317] compared the zero-shot performance of ChatGPT on the SOLD dataset against a number of other Sinhala models. They conclude that Zero-shot ChatGPT perform better in Sinhala hate speech detection than models that were not fine-tuned on SOLD. Dmonte et al. [141] used the SOLD data set to verify the new MT-Offense124124124https://github.com/LanguageTechnologyLab/MT-Offense data set they created by using NMT to translate the English offensive language benchmark dataset, OLID [142]. They report the best F1 of 0.550 for the translation done using nllb-200-3.3B [134].
@@ 3.14 Word Sense Disambiguation
There have been multiple attempts to do word sense disambiguation (WSD) [322, 323, 324, 325, 326] for Sinhala. For this, Arukgoda et al. [191] have proposed a system named Aruth based on the Lesk Algorithm[327]. An online tool125125125http://aruth.herokuapp.com/, an API126126126https://bit.ly/3sJEYbS of the algorithm, and code along with data on github127127127https://github.com/jseanm1/aruthSWSD are available. For the same task, Marasinghe et al. [328] have proposed a system based on probabilistic modeling. A dialogue act recognition system which utilizes simple classification algorithms has been proposed by Palihakkara et al. [329]. A word sense disambiguation tool named Sinsense was introduced by Subasingha [330]. They used cross-lingual sense disambiguation where English sense disambiguation was used to obtain Sinhala sense disambiguation. However, neither their tool nor their full research is publicly available.
@@ 3.15 Text Summarizing
A deterministic process flow for automatic Sinhala text summarizing was proposed by Welgama [331]. The study by Wimalasuriya [332], which has the same name as the above work by Welgama [331], uses the graph-based TextRank algorithm for automatic Sinhala text summarizing. The use case of Sinhala Text summarization for government gazettes was explored by Jayawardane [333]. The study by Rathnayake et al. [334] compared the results of extractive and abstractive summarization on Sinhala textbooks. The study by Jahan and Wijesekara [335] compared the abstractive summarizing methods of TF-IDF and Text-Rank for Sinhala using ROUGE as the evaluation score. They concluded TF-IDF to be the superior choice. Hasan et al. [151] introduced the XL-sum128128128https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website129129129https://www.bbc.com/sinhala. The M3LS130130130https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al. [152] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website131131131https://www.bbc.com/sinhala. Further, they claim that MT5 [336] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set. The first Sinhala multi-document summarizing dataset, M2DS132132132https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al. [153]. The M2DS dataset consists of 23.5k Sinhala documents in 5.5k clusters with each cluster having a golden summary. Thier code and pre-trained models are also available on github133133133https://github.com/KushanMH/m2ds.
@@ 3.17 Phonological Tools
On the case of phonological layer, a report on Sinhala phonetics and phonology was published by Wasala and Gamage [338]Wickramasinghe et al. [339] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems. The Massively Multilingual Speech (MMS)134134134https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al. [145] has Sinhala data for the Spoken Language Identification (LID) task. The meta-study by Al-Fraihat et al. [340] compared the status of Sinhala speech recognition research against 17 other languages.
@@ 3.17.1 Text-to-Speech
Based on the earlier work by Weerasinghe et al. [341]Wasala et al. [342] have developed methods for Sinhala grapheme-to-phoneme conversion along with a set of rules for schwa epenthesis. This work was then extended by Nadungodage et al. [343]Weerasinghe et al. [344] developed a Sinhala text-to-speech system. However, it is not publicly accessible. They internally extended it to create a system capable of helping a mute person achieve synthesized real-time interactive voice communication in Sinhala [345]. A rule based approach for automatic segmentation of a small set of Sinhala text into syllables was proposed by Kumara et al. [346]. An ew prosodic phrasing method to help with Sinhala Text-to-Speech process was proposed by Bandara et al. [347, 348, 349]Sodimana et al. [350] proposed a text normalization methodology for Sinhala text-to-speech systems. Further, Sodimana et al. [351] formalized a step-by-step process for building text-to-speech voices for Sinhala. Both Jayamanna [352] and Mishangi [353] have created Sinhala document readers for visually impaired persons to be used on Android devices. An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al. [354]. The works by Lakmal et al. [355] and Senarathna et al. [356] adapted MaryTTS [357] to synthesize Sinhala speech. The study by Jayawardhana et al. [358] used Deep Voice [359] for Sinhala and English TTS. Gamage et al. [360] included a Sinhala text-to-speech module as one of the three modules present in their currency recognition system. The study by Madhusha et al. [361] claims to have created a mobile app with Sinhala Text-to-Speech and OCR to read books for visually impaired students. Praveen et al. [362] claim to have created a Sinhala TTS system with a custom neural network that archives 98% accuracy.  Anuradha and Thelijjagoda [363] proposed a machine translation system to convert Sinhala and English Braille documents into voice. A separate group has done work on Sinhala text-to-speech systems independent to above [364].
@@ 3.17.2 Speech-to-Text
On the converse, Nadungodage et al. [365] have done a series of work on Sinhala speech recognition with special notice given to Sinhala being a resource poor language. This project divides its focus on: continuity [366], active learning [367], and speaker adaptation [368]. A Sinhala speech recognition for voice dialing which is speaker independent was proposed by Amarasingha and Gamini [369] and on the other end, a Sinhala speech recognition methodology for interactive voice response systems, which are accessed through mobile phones was proposed by Manamperi et al. [370]. A Sinhala speech to Unicode text converter for the disaster relief domain was proposed by Prasangini and Nagahamulla [371]Priyadarshani [372] proposes a method for speaker dependant speech recognition based on their previous work on: dynamic time warping for recognizing isolated Sinhala words [373], genetic algorithms [374], and syllable segmentation method utilizing acoustic envelopes [375]. The method proposed by Gunasekara and Meegama [376] utilizes an HMM model for Sinhala speech-to-text. A Sinhala speech recognizer supporting bi-directional conversion between Unicode Sinhala and phonetic English was proposed by Punchimudiyanse and Meegama [377]. The work by Karunanayake et al. [378] transfer learns CNNs for transcribing free-form Sinhala and Tamil speech data sets for the purpose of classification. Dilshan [379] conducted a study for the specific use case of transcribing number sequences in continuous Sinhala speech.  Gamage et al. [380] explored the use of combinational acoustic models such as Deep Neural Network - Hidden Markov Model (DNN-HMM) [381] and Subspace Gaussian Mixture Model (SGMM) [381] in Sinhala speech recognition. In the follow-up work, Gamage et al. [382] extended that work with end-to-end Lattice-Free Maximum Mutual Information (e2e LF-MMI) model [383] which is claimed to be a viable solution for low resource language speech recognition by Carmantini et al. [384]. However, it was shown that the new model slightly underperforms compared to the state-of-the-art result. Later, they conducted further development on their model in a follow-up work [385].  Karunathilaka et al. [386] explore Sinhala speech recognition using deep learning models such as: pre-trained DNN, DNN, TDNN, TDNN+LSTM. The first half of the study by Arafath [387] dealt with recognizing Sinhala speech using LSTMs. Gamage et al. [360] included a Sinhala speech recognition module as one of the three modules present in their currency recognition system. Time-delay neural network architectures (including multistream CNN architecture) were used for acoustic modeling of Sinhala Automatic Speech Recognition (ASR) by Warusawithana et al. [388]. They have used the Kaldi speech recognition toolkit [389] for training the ASR models. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [390], proposed a Sinhala speech recognition system. The study by Azir et al. [391] attempts to identify number sequences spoken in Sinhala. TacoSi introduced by Arachchige and Weerasinghe [392] is based on Tacotron [393] and has been evaluated with 10 human evaluators to determine its text-to-speech quality. Nanayakkara [394], Nanayakkara and Weerasinghe [395] used DeepSpeech135135135https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition. Gunarathne et al. [396] used an earlier version of the CMUSphinx toolkit136136136https://cmusphinx.github.io/ to transcribe Sinhala speech to text. The later work by Akesh and Meegama [397] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [398, 399] to automatically generate Sinhala subtitles from Speech. The work by Wickramaarachchi et al. [400] also used MFCC features but focused on Sinhala speech intonation recognition for the purpose of detecting speech impediments in young children. Dissanayaka et al. [401] have introduced Word Sri an application that is voice-activated and capable of Sinhala grammar checking and plagiarism checking. tha [402] have introduced a Sinhala text to Speech dataset137137137https://github.com/SiTa-SpeakerDiarization/SiTa transcribed from 60 videos with a total of 602 minutes. As the baseline, they report the lowest Diarization Error Rate (DER) of 6.0 for Powerset Cross-Entropy Diarization [403]. While the data set link points to a github repository, it does not contain the data set; it points to a project website138138138https://sita-speakerdiarization.github.io/ which then points to a download page139139139https://rtuthaya.staff.uom.lk/resources/dataset/44 where you need to send a request to the authors to obtain the data set.